What is the maximum amount in moles of P2O5 that can theoretically be made from 112 g of P4 and excess oxygen?
1
Expert's answer
2017-11-08T11:45:08-0500
4P + 5 O2 = 2 P2O5
n (P) = m / M = 112 g / 30.97 (g/mol) = 3.616 mol
n (P2O5) = 0.5 * n (P) = 3.616 mol * 0.5 = 1.808 mol
Answer:
1.808 mol